In a parallelogram ABCD, P divides AB in the ratio 2 : 5 and Q divides DC in the ratio 3 : 2. If AC and PQ intersect at R, find the ratios AR : RC and PR : RQ
1
Expert's answer
2019-09-17T11:28:57-0400

"AK||PQ" and "PQ||LC"

"AO=h" (height)

"APR" and "CQR" similar triangles "AR\/RC=PR\/RQ"

"AP=2*X\\
PB=5*X\\
DQ=3*Y\\
QC=2*Y\\
2*X+5*X=3*Y+2*Y\\
X/Y=5/7\\
Area APQK=2*X*h\\
Area PLCQ=2*Y*h\\
AreaAPQK/AreaPLCQ=AR/RC=PR/RQ=(2*X*h)/(2*Y*h)=X/Y=5/7"

Answer: "AR\/RC=PR\/RQ=5\/7"